OpenStudy (anonymous):

Julian deposits money into two different savings accounts. He deposits $600 more into Account B than Account A. After a certain amount of time, Account A has earned $180 at a rate of 4%, and Account B has earned $375 at a rate of 5%. How much did Julian initially deposit into each account? What equation is used to solve this problem? What does each variable represent? Identify the given information. What do you know about P? Enter the given information into the equation. Solve the equation. State the solution.

OpenStudy (anonymous):

This is apart of my online class. For part B of my assignment, I had to make a similar problem, which is this: A) Ashley deposited money into two different savings accounts. She deposited $10,000 more into Account B than Account A. After a certain amount of time Account A earned $1,000 with a rate of 10%, and Account B earned $3,000 with a rate of 15%. How much did Ashley individually deposit to each account? B) C) Account A - $10,000 Account B - $20,000 But I don't know how to solve it with and equation. I only know the answer because I made it up. Please help!

OpenStudy (anonymous):

here is the equation for compounding for a single initial sum \[FV=PV(1+i)^n\] Account A would look like this: PV_a =PV_a FV_a = PV_a+180 i_a = 4% \[PV_a+180 = (PV_a)(1+0.04)^n\] Account B: PV_b = PV_a+600 FV_b = 375+PV_b i_b = 5% \[375+PV_a+600 = (PV_a+600)(1+0.05)^n\] solve for PV_a and n

OpenStudy (anonymous):

thank you!